OpenStudy (anonymous):

solve for x and y {y=-2x+1 {3x-2y=9

OpenStudy (anonymous):

You can solve by graphing, elimination or substitution. Which method do you want me to use out of the last 2 options?

OpenStudy (anonymous):

put the value of y from first to 2nd one-------this is called substitution method 3x-2(-2x+1)=9 3x+4x-2=9 7x-2=9 7x=9+2 7x=11 x=11/7 put in any equation u will get y:)

OpenStudy (anonymous):

substitution i can find x but I can't find y I'm not sure how to do it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Elimination: Put the -2x in the first term on the left hand side by adding 2x 2x+y=1 (1) 3x-2y=9 (2) 2x+y=1 multiply by (2) you get: 4x+2y=2 3x-2y=9 Add together: 7x=11 x = 11/7 Put x=11/7 in any of the original equations and solve for y: 2*(11/7)+y=1 solve for y = 1-22/7 y=-15/7

OpenStudy (anonymous):